### v1

# Uncovering the role of semantic and acoustic cues in normal and dichotic listening

Abstract
Despite extensive research, the precise role of acoustic and semantic cues in complex speech perception tasks remains unclear. In this study, we propose a paradigm to understand the encoding of these cues in electroencephalogram (EEG) data, using match-mismatch (MM) classification task. The MM task involves determining whether the stimulus and response correspond to each other or not. We design a multi-modal sequence model, based on long short term memory (LSTM) architecture, to perform the MM task. The model is input with acoustic stimulus (derived from the speech envelope), semantic stimulus (derived from textual representations of the speech content), and neural response (derived from the EEG data). Our experiments are performed on two separate conditions, i) natural passive listening condition and, ii) an auditory attention based dichotic listening condition. Using the MM task as the analysis framework, we observe that - a) speech perception is fragmented based on word boundaries, b) acoustic and semantic cues offer similar levels of MM task performance in natural listening conditions, and c) semantic cues offer significantly improved MM classification over acoustic cues in dichotic listening task. Further, the study provides evidence of right ear advantage in dichotic listening conditions.
Index Terms:
EEG, Match-mismatch classification, multi-modal modeling, Speech perception, acoustics and semantics.I Introduction
Humans speech perception is a complex cognitive process involving the integration of acoustic and contextual cues. The speech envelope is the most commonly used acoustic feature for analyzing neural entrainment and is shown to be synchronized with neural oscillations in the auditory cortex [1]. Furthermore, recent research [2] suggests that neural entrainment to speech is not limited to the speech envelope as it also extends to other acoustic features such as phonemic content. However, a systematic understanding of the speech attributes that are encoded in the brain is somewhat lacking. In this paper, we investigate the importance of word boundary information and relative role of semantics and acoustics in parsing the input speech under normal and dichotic listening conditions.
Continuous natural speech stimuli evoke a brain response that can be detected with neuroimaging techniques like electroencephalography (EEG) or magnetoencephalography (MEG). Electroencephalography (EEG) recording of brain responses is one of the simplest approaches to understand the neural processes involved in speech comprehension. The EEG is a non-invasive neural imaging technique that measures electrical activity in the brain by placing electrodes on the scalp [3]. Multiple studies have demonstrated that EEG, recorded during speech listening tasks, contains information about the speech stimuli [1, 4, 5]. Liberto et al. [1] demonstrates that EEG reflects categorical processing of phonemes in continuous speech. Additionally, Broderick et al. [4] showed that EEG captures linguistic representations such as semantic dissimilarity. Similarly, Soman et al. [5] revealed that the EEG recorded during the listening state carries language information of the stimulus.
The design of studies intended to decode the role of different features during speech perception has been an important research area in auditory neuroscience. Early studies [6, 7, 8, 9] in this area explored linear models to establish a relationship between continuous natural speech and EEG responses, classified as forward, backward, or hybrid models. These models predict either the EEG from speech stimuli or reconstruct the speech from EEG responses. The correlation between the predicted and the reference signals has been used as a measure of neural tracking in numerous studies [10]. However, due to the non-linear nature of the auditory system [11], linear models may not be sufficient for capturing the stimulus-response interaction during speech perception. Deep neural networks have recently been used to compare and analyze EEG responses for spoken stimuli [12, 13, 14, 15].
The match-mismatch (MM) tasks have been proposed as an alternative task to the previously explored forward/backward models [14, 16]. In this case, the goal is to determine if a segment of brain response (like EEG) is aligned with a segment of spoken stimulus. This approach, measured by MM classification accuracy, provides a simpler and more interpretable metric compared to the traditionally employed correlation metrics. The correlation approaches often suffer from low correlation values (typically around -%), limiting their applicability [10]. Additionally, correlation metrics are more susceptible to the influence of outliers compared to a classification model. On the other hand, the match-mismatch task, framed as a binary classification problem, is considered an easier task for models to learn when compared to the more complex regression tasks performed by forward-backward models.
In most of the previous studies using the MM task, the auditory stimulus of a fixed duration (s) is processed through a series of convolutional and recurrent layers [12, 15, 16]. The recent study by Soman et al. [17] proposed a deep learning model to investigate stimulus-response alignment at the sentence level. This work showed the importance of incorporating word-level segmentation information for auditory stimulus-response modelling. The work improved the MM classification accuracy compared to the previous works [12] during natural speech listening.
In addition to analysing speech perception during natural speech listening, we also aim to analyse how the brain comprehends speech in complex listening conditions, such as in dichotic listening. Such auditory attention decoding (AAD) experiments have been used to investigate auditory processing and attention mechanisms in the brain [6, 18]. Auditory attention decoding refers to the process of decoding or identifying the specific auditory stimulus or sound that an individual is paying attention to when presented with a mixture of auditory stimuli. One such complex listening condition is the dichotic listening, in which participants are presented with different auditory stimuli simultaneously in each ear. This capability of the human brain is termed ”The Cocktail party effect”, which refers to the ability of the human brain to selectively focus attention on a specific auditory stimulus while filtering out other competing sounds in a noisy environment. Our study aims to identify the specific cues to which the subject directs their attention during this task. Additionally, we aim to analyze the influence of word boundary information on auditory attention decoding (AAD).
In this study, we propose the MM framework for auditory stimulus-response (EEG) modeling with a sentence-level model that incorporates word segmentation information [17]. This framework is designed to capture the neural responses in both mono-aural and dichotic listening scenarios. For dichotic listening, we formulate MM task as a means to perform auditory attention decoding (AAD), aiming to identify the channel the subject was attending. To conduct this study, we utilize two distinct speech-EEG datasets [4]: one collected during a natural speech listening task and the other during a dichotic listening task.
Our approach involves employing a deep learning network for stimulus-response modeling, which consists of three separate sub-networks: one for processing the EEG signals, one for processing the acoustic features of the stimulus and the other for processing the semantic features of the stimulus. The EEG sub-network comprises a series of convolutional layers, followed by word boundary-based average pooling, and a recurrent layer that incorporates inter-word context. The stimuli sub-network is configured based on the input features.
The acoustic feature used to represent phonological traits is the speech envelope. Additionally, the semantic feature employed is the word2vec [19], which represents words as vectors in a high-dimensional space based on their contextual usage. We conduct a comparative analysis to evaluate the relative contributions of acoustic and semantic features in both natural and dichotic listening conditions. The results indicate that semantic features exhibit higher accuracy in match-mismatch (MM) classification during dichotic listening, while both features perform similarly in natural listening conditions.
This paper extends the preliminary work [17] in multiple ways. The major contributions of this paper are:
Formulate a paradigm for auditory stimulus-response (EEG) modeling through a match-mismatch (MM) task in natural listening as well as in dichotic listening.
In the mono-aural natural speech listening task, the MM task performance of audio signals is enriched by the induction of word boundaries. Further, the textual information also provides statistically similar MM performance compared to the acoustic features.
In the dichotic speech listening task, the MM task performance of textual features is higher than that of the audio signal, indicating that EEG signals encode higher-level semantic information which may be accessed to separate the acoustic channels and pay selective attention.
For both these tasks, we propose a multi-modal sequence model based on LSTM architecture. The performance of the multi-modal model significantly improves over the individual modality of speech, indicating that the EEG signal jointly encodes the semantic and acoustic content of the stimulus. Further, we also illustrate that the proposal significantly improves over previously proposed models for MM task [12].
For dichotic listening, we show evidence of right ear advantage through the MM classification performance.
Impact of the proposed study
The results of these experiments find applications in robotics [20, 21], brain-computer interface (BCI), and hearing aid applications [22]. Hearing loss and the associated use of hearing aids are on the rise in the general population [23]. The performance of hearing aids is significantly impacted in situations where background noise involves multiple talkers [24], known as the cocktail party problem. Addressing the challenge of assisted listening in multi-talker environments could have broad societal benefits in terms of communication and mental health [25]. AAD leverages knowledge of the listener’s auditory intent (attention) to isolate and enhance the desired audio stream, while suppressing others [26, 27, 28].
Moreover, auditory attention detection algorithms could be integrated with passive Brain-Computer Interface (pBCI) systems for applications in education and interactive music performances [29]. An early attempt in this direction was made by Cho et al. [30], who developed an attention enhancement system for ADHD children using EEG biofeedback in a virtual classroom environment. The AAD method suggested in this work can be used to continuously track the attended auditory source and provide feedback to the user [29].
II Methods
II-A Dataset
We experiment with two publicly available speech-EEG data sets111https://doi.org/10.5061/dryad.070jc released by Broderick et al. [4]. These scenarios include subjects listening to uninterrupted, natural speech and a situation where the subjects are exposed to the cocktail party effect. In the latter case, the subjects listen to two distinct audio streams simultaneously, each presented in a separate ear. For the subsequent discussion, we refer to the first dataset as the natural speech dataset and the second one as the dichotic (termed as ”cocktail party” in the original dataset) dataset.
In both experiments, the stimuli were presented using Sennheiser HD650 headphones with the software provided by Neurobehavioral Systems (http://www.neurobs.com). The participants were instructed to keep their gaze fixed on a crosshair displayed on the screen throughout each trial and to minimize activities such as eye blinking and other motor movements.
II-A1 Natural Speech Dataset
The natural speech dataset contains electroencephalographic (EEG) data recorded from subjects as they listened to continuous speech. The subjects listened to a professional audio-book narration by a single male speaker of a well-known work of fiction. For each subject, the data consists of trials of similar length, each containing 180s of audio. The trials preserved the chronology of the storyline without repetitions or breaks. The sentence start and end time and the word-level annotation of the speech recordings are provided. The word segmentation is obtained using a speech recognition-based aligner [31]. The EEG data were acquired using -channel BioSemi system at a sampling rate of Hz, while the audio data is played at kHz. Overall, the speech-EEG data amounted to a duration of hours.
II-A2 Dichotic Dataset
The dichotic dataset used in this study comprises EEG recordings obtained from subjects. Each participant listened to a total of trials, each lasting seconds. They were presented with two well-known fictional works, where one story was played to the left ear and the other to the right ear. Different male speakers articulated each story. The participants were divided into groups of and individuals (with one subject excluded). Each group of listeners was instructed to focus their attention solely on the story presented in either the left or right ear for all the trials. Following each trial, the participants were required to answer multiple-choice questions about the stories played in left/right ear, each offering four possible choices. To maintain consistency, the audio streams of each story within a trial were normalized to have the same root mean squared (RMS) intensity. To prevent the unattended story from capturing the participants’ attention during silent periods in the attended story, any silent gap exceeding seconds was truncated.
II-B EEG Pre-processing
The EEG pre-processing pipeline utilized in this study for both datasets was based on the CNSP Workshop 2021 guidelines222https://cnspworkshop.net/resources.html and implemented using the EEGLAB software [32]. First, a low-pass Butterworth filter with a cutoff frequency of Hz was applied to the EEG signal. This filter helps attenuate high-frequency noise and artifacts that are unrelated to the neural activity of interest. Next, a high-pass Butterworth filter with a cutoff frequency of Hz was employed to remove DC offsets or drifts in the signal. Following the filtering steps, the EEG data were down-sampled to Hz.
A channel rejection method based on the EEG data’s variance was employed to identify and replace noisy channels. The channels with excessively high variance, indicating potential artifacts or poor signal quality, were considered unreliable and replaced using a spline interpolation technique. This interpolation is performed using the remaining channels to preserve the spatial information of the EEG data.
After channel rejection and interpolation, the EEG channels were re-referenced to a specific set of external channels known as mastoids. This re-referencing step helps minimize the effects of common noise sources and improves the interpretability of the EEG data by providing a reference point.
Lastly, to normalize the data and ensure consistency across channels and trials, a z-score transformation was applied. This process computes the z-score independently for each channel and each trial, by subtracting the mean followed by normalization with standard deviation.
II-C Match-mismatch classification task for natural listening
Our work uses the MM task as the stimulus-response model to study the neural entrainment. Figure 1 illustrates this paradigm in detail. The classification model is trained to maximize the similarity between the EEG and the “attended” speech segment while minimizing the similarity between the same EEG data and the “un-attended” audio. In this study, the speech segment is considered at a sentence-level. We also compare with prior works [12, 33], which perform this task at the sentence level. In the natural listening condition, the stimulus corresponding to the EEG response segment is the matched speech while speech from another sentence in the same session is chosen as the mismatched speech. Selecting mismatched samples from the same trial makes the classification task challenging enough to encourage the model to learn the stimulus-response relationships. This sampling approach also avoids the chance of memorizing the speech features along with their labels. The details of MM task for dichotic listening setup are given in the following sub-section.
II-D Auditory Attention Decoding as a Match-mismatch task
Figure 1 illustrates this paradigm in detail. The segment of speech that the subject is instructed to attend is considered as the match segment, while the segment of speech played in the other ear is considered as the mismatch segment. This task is more challenging than MM tasks on the natural speech dataset. The behavioral experiment that followed the listening experiment showed that the subjects were able to comprehend and understand the segment they attended to while having difficulty answering questions about the unattended speech. We devise a trial-selection based on behavioral scores (Figure 3), which is described in Sec. II-H.
II-E Acoustic Feature - Speech Envelope
The speech envelope represents the variations in the amplitude of the speech signal over time. It is obtained by extracting the magnitude of the signal’s analytic representation, using the Hilbert transform. The temporal envelopes of sounds contain critical information for speech perception [34, 35]. It has been shown that auditory cortex can temporally track the acoustic envelopes [2]. The strength of cortical envelope tracking may serve as an indicator of the extent of speech perception ([36, 37, 9]). Speech envelope provides valuable information about the overall shape and dynamics of the speech signal, including prosodic features, syllabic structure, and phonemic transitions. However, it should be noted that the speech envelope does not contain information regarding the semantics of the speech input.
II-F Semantic Feature - Word2vec text embedding
The text transcription of the speech stimulus is provided in the dataset. This study used this text data to obtain features representing the semantics of the stimulus speech signal. Semantic vectors for content words were derived using the word2vec algorithm [19].
Word2Vec (w2v) is a popular word embedding technique proposed by Mikolov et al. [19]. This algorithm generates a vector representation for each word. This study used pre-trained vectors trained on a subset of the Google News dataset (about 100 billion words). The model contains 300-dimensional vectors for 3 million words and phrases. The embeddings were obtained using a data-driven approach outlined in [38]. These pre-computed word vectors are publicly available333https://code.google.com/archive/p/word2vec/. The fundamental notion of w2v embedding is that words with similar semantics tend to be closer to each other in their vector space representation. In our work, the text embeddings were computed for each word in the sentence. The w2v vectors of constituent words were concatenated together to obtain the sentence representation. This representation was fed as input to the models described in the following sections.
II-G Model architecture
II-G1 Acoustic Encoding
The speech signal representation is the speech envelope of dimension , where denotes the duration of a speech sentence at Hz. Similarly, the EEG data for the same sentence is denoted as , which is of dimension .
The EEG sub-network (top row of Figure 2) consists of a series of convolutional layers and LSTM layers. The convolutional layers implement -D and -D convolutions with and kernel sizes, respectively. Further, the -D CNN layers also introduce a stride of to further down-sample the feature maps. The speech sub-network (middle row in Figure 2) has a 2D convolutional layer of kernel size 16 and stride (1,3). For acoustic encoding, we utilize both the speech and EEG subnetworks in a parallel neural pipeline, as depicted in Figure 2, without any weight sharing.
The word boundary information available in the dataset is provided at the equivalent sampling rate (both EEG and audio representations at Hz). The audio and EEG feature maps are average pooled at the word level using the word boundary information. As a result, for a given sentence, the EEG and speech branches generate vector representations sampled at the word level. An LSTM layer models the inter-word context from these representations. This layer is included in both the stimulus (speech) and response (EEG) pathways. The last hidden state of the LSTM layer, of dimension , is used as the embedding for the stimulus/response, denoted as /, respectively.
We propose the Manhattan distance [39] between the stimulus and response embeddings. The similarity score is computed as,
The similarity score for the matched pair () and mismatched pair () are computed. The model, with a dropout factor of , is trained using a binary cross-entropy loss, with [, ] mapped to [, ] targets.
II-G2 Semantic Encoding
To compare the encoding of text features with speech features, we used a text-EEG match mismatch classification task. We used the text-EEG model sub-network (bottom row of Figure 2) to learn the mapping between the w2v embeddings and the EEG response. The EEG sub-network resembles the acoustic encoding model, while the text sub-network consists of two-layers of LSTM architecture with w2v embeddings at word-level as input features. The convolution layers were not used as the word2vec (w2v) features do not contain 2-D information. The similarity scoring component is identical to the acoustic encoding model (Eq. 1).
II-G3 Joint Encoding of Acoustics and Semantics
In addition to exploring the effect of acoustic and semantic features individually, we perform an investigation of the joint training of the MM model with both the modalities. The sub-networks are combined by defining a joint-loss function over the two modalities (detailed in Sec. II-I). For a given match segment, the EEG response corresponds to the attended audio/text features, while in the mis-matched example, the audio/text features are chosen from a different region (natural speech) or from the stream played in the other ear (dichotic listening).
II-H Trial selection in Dichotic listening
As part of the dichotic listening task, participants were required to answer multiple-choice questions after each trial to assess their comprehension and, consequently, their attentiveness to the played stories [4]. They were asked questions each from both the stories. The scores obtained were normalized to a range of to . For all the experiments that we report in this work, we include trials in which the participants achieved a comprehension score of % or more for attended audio, while achieving a score of % or less for the un-attended audio. This choice of scores for attended / un-attended streams ensured that the “matched” / “mis-matched” EEG-audio/text pairs had a clear distinction. The selection is elicited in Figure 3. While this selection led to a removal of % of the total trials, the selection ensured a reliable label for the training and evaluation.
II-H1 Training and Evaluation Setup
The experiments are run with a batch size of . The models are trained using Adam optimizer with a learning rate of and weight decay parameter of . The models are learned with a binary cross-entropy loss.
Overall, the natural speech dataset contains stimulus trials, while the dichotic listening dataset contains trials. A subset of trials from all subjects constituted the training set, while the remaining unseen trials of all subjects served as the test data. In each validation fold, we randomly chose trials to form the test data. This resulted in -fold cross-validation for the natural speech dataset and -fold cross-validation for the cocktail party dataset. The average classification accuracy across validation folds is reported in the results section.
To assess statistical significance, we employed the Wilcoxon signed-rank test [40]. This is performed in a paired setting, where the predicted probabilities from the two models are used in the comparison. We have used an level of for determining the statistical significance. For the unpaired t-test, we have also performed the Wilcoxon rank-sum test.
II-I Multimodal Loss Function and Joint training
Similarity | |||||||
measure | 0 | 0.2 | 0.4 | 0.5 | 0.6 | 0.8 | 1 |
82.00 | 82.12 | 82.87 | 83.76 | 82.92 | 79.46 | 77.45 | |
81.20 | 81.60 | 80.00 | 78.46 | 78.00 | 77.60 | 75.71 | |
77.76 | 79.00 | 78.72 | 78.12 | 77.57 | 76.97 | 74.61 |
A hyper parameter is introduced, which controls the contribution of the text and speech modalities in the joint training. We explore three loss function choices which combine the modalities,
Here, implies that only the semantic network (text-EEG) is considered, implies that only the acoustic network (speech-EEG) is considered while implies joint modeling (EEG-speech+text) with equal weight to speech and text modalities. During training, the speech, text and EEG data are loaded and the value of is randomly selected from . In this way, a single model is trained with all three combinations, of text alone (), speech alone () and speech-text multimodal (). During inference, the model can be input with EEG-text samples (), EEG-speech samples () as well EEG-speech+text (). This flexibility helps us understand the role of acoustics and semantics separately from the same model.
III Results
III-A Choice of loss function
For dichotic listening condition, we experiment with different values during inference. As shown in Table I, it is observed that semantic encoding in EEG achieves a higher MM performance than the acoustic encoding network. Further, the similarity measure proved to be best performing similarity measure. The best MM classification results are achieved with . All the subsequent experiments use this configuration with similarity measure.
III-B Impact of Acoustic and Semantic Cues
cells = c,
hlines,
vline2 = -,
hline1,4,6 = -0.08em,
hline5 = -dashed,
hline3 = -dashed,
colsep = 4pt
Listening
Condition & Speech
envelope
(=1) Text
word2vec
(=0) Multimodal
(Acoustic
+Semantic)
(=0.5)
Natural (Baseline) [12] 65.23 – –
Dichotic (Baseline) [33] 62.00 – –
Natural 93.63 93.24 93.38
Dichotic 77.45 82.00 83.76
Table II presents the results of the MM classification accuracy between speech stimuli and their corresponding EEG responses, focusing on different listening conditions. Compared to the MM classification of natural speech under normal listening conditions, the task of auditory attention decoding (AAD) in dichotic listening is more challenging.
For natural speech conditions, both semantic and acoustic features result in similar accuracy (Wilcoxon signed rank test, ). For dichotic listening, word2vec features result in significant improvement in accuracy compared to speech envelope features (Wilcoxon signed rank test, ). The scatter plot for prediction probabilites for dichotic listening condition and for a fold in validation is shown in Figure 4.
III-B1 Importance of word boundary information
Word-level segmentation holds significant importance in speech perception during normal listening conditions [17]. To investigate the role of word segmentation in auditory attention decoding, we conducted a comparison of match-mismatch classification performance using the proposed model with and without word boundary information in the dichotic listening condition. The word boundary segmentation is used only for the EEG and acoustic features, while the word2vec features are inherently sampled at word-level.
As seen in Table III, word level segmentation plays an important role in auditory attention decoding. Word boundary information has impact on both semantic features (Wilcoxon rank-sum test, ) and envelope features (Wilcoxon rank-sum test, ). Word boundary information also shows a significant impact on the joint speech-text model.
column2-3 = c,
hlines,
hline3-4 = -dashed,
vline2-3 = -
Stimulus Feature & Without
word-boundary With
word-boundary
Envelope (=1) 67.09 77.45
word2vec (=0) 77.35 82.00
Multi-modal (=0.5) 76.69 83.76
III-B2 Importance of accurate word boundaries
We conducted another ablation study to understand the impact of accurate word boundary information. In this setting, the model is fed with random word boundaries with varying granularity. Each sentence is segmented to a fixed number of regions, and their boundaries are chosen at random. The results are reported in Table IV. The performance does not significantly improve with acoustic feature (Wilcoxon rank-sum test, ), semantic feature (Wilcoxon rank-sum test, ), or with multi-modality (Wilcoxon rank-sum test, ) as we increase the number of random word boundaries. However, the use of accurate word-boundaries is significantly better in all cases over the ones with random word boundaries.
row1 = l,
cell11 = l,
cell22-6 = c,
cell32-6 = c,
cell42-6 = c,
hlines,
vline2 = -,
vline6 = -dashed,
hline3-4 = -dashed,
Stimulus
Feature & 2 3 4 5 True word
boundary
Envelope
(=1)
66.13 66.16 67.39 70.89 77.45
word2vec
(=0)
78.95 76.29 78.58 78.19 82.00
Multi-modal
(=0.5)
79.11 79.20 79.55 79.41 83.76
In the second set of ablation experiments, we provide accurate word boundary information but skip the word boundary information at every -th word. These results are reported in Table V. For example, ‘Skip-3’ in this table means that we skip the word boundary value at every -rd location from the word boundary value list of a given sentence. The pooling is done with the rest of the available word boundaries for these experiments. A higher value of “n” in “Skip-n” indicates an increased use of word boundary information.
column2-4 = c,
hlines,
vline2 = -,
hline3-4 = -dashed,
Stimulus & Skip-2 Skip-3 Skip-4 Skip-4 No Skip
Env. (=1) 74.94 79.10 75.49 80.20 77.45
w2v (=0) 81.1882.78 78.35 80.93 82.00
M-modal 82.69 84.18 80.21 83.47 83.76
III-B3 Region wise Analysis
cells = c,
hlines,
vline2 = -,
hline1 = -0.08em,
hline3,4,5,6 = -dashed,
Brain
Region & Speech
envelope
(=1) Text
word2vec
(=0) Multimodal
(Acoustic
+Semantic)
(=0.5)
Frontal 69.70 74.85 75.65
Central 57.39 62.30 62.47
Parietal 71.09 78.03 78.57
Temporal 77.00 79.33 81.14
Occipital 73.33 79.00 79.98
We performed an experiment to understand which regions of the brain were relatively more involved and influential in performing the attention task in a dichotic listening setting. While spatial localization in EEG is not accurate, the analysis may indicate potential regions which carry more information for the MM task.
The electrodes of the -electrode EEG setup are segregated into five regions namely frontal, central, parietal, temporal and occipital. We use the - map to perform this segregation. Only the subset of electrodes belonging to a region are used in training and cross-validation. As shown in Table VI, the temporal region performs significantly better in MM classification in comparison with other regions (Wilcoxon rank-sum test, ).
III-B4 Right Ear Advantage
cells = c,
hlines,
vline2 = -,
hline1,4 = -0.08em,
hline3 = -dashed,
Stimulus & Speech
envelope
(=1) Text
word2vec
(=0) Multimodal
(Acoustic
+Semantic)
(=0.5)
S1 (left ear) 74.14 79.12 83.52
S2 (right ear) 79.55 84.75 87.37
We conducted a set of ablation experiments aimed at identifying which stimulus group (attention to left versus right ear) elicited the strongest performance in the MM classification task. As mentioned in Sec. II-H, the listeners are instructed to attend to stories played in left or right ear. Thus, we split the data into two parts, based on which side of the story the participant was instructed to attend, and chose the trials which positively reflected this instruction in the behavioral data. An initial observation was that the trial selection (discussed in Sec. II-H) retained more trials ( %) when the subjects were asked to attend to the right ear than when they were asked to attend to the left ear (%). The MM task results from the two subject splits are shown in Table VII after balancing the training data of each group to the same number of samples.
We observe that the subject group instructed to listen to the right ear (S2) had significantly improved MM task performance than the subject group (S1) instructed to attend to the left ear (Wilcoxon rank-sum test ). This improvement was consistently seen on all modalities (speech, text and multimodal). We find that this result further aligns with several established studies indicating a right ear advantage in dichotic listening (For example, early work reported in Kimura et al. [41] and more recent studies in Payne et al. [42])
IV Discussion
IV-A Word Boundary Information and Impact on MM Task
With a set of experiments, we have attempted to validate the hypothesis that speech perception in the brain is segmented at the word level. The incorporation of word boundary information yields statistically significant improvements compared to the baseline model, demonstrating the impact of this information in the neural tracking of speech. Both normal and dichotic listening tasks benefit from the use of accurate word information. Further, random choice of word-boundaries do not result in significant improvements.
Our findings suggest that modelling EEG responses and stimuli yield superior performance during natural speech perception. Furthermore, we observed that the relative importance of textual and acoustic content is approximately similar (Wilcoxon signed rank test, ) in the natural listening condition. However, during dichotic listening, the human brain exhibits a prioritization of perceiving the semantics over the acoustic characteristics of the attended speech (Wilcoxon signed rank test, ). The results indicate that word boundary information plays a crucial role in speech perception during the dichotic listening task. Nonetheless, it’s worth noting that recognizing all boundaries during dichotic listening is not as critical compared to the natural speech listening condition, as there is no significant drop in accuracy (Wilcoxon rank-sum test, ) between skip-2 and skip-3 (as well as its subsequent instances)as shown in Table V. This result supports the notion that the human brain prioritizes comprehending the content over phonological aspects of the stimulus during challenging listening conditions, and this comprehension likely occurs at the phrase level or higher.
IV-B Natural Speech: Speech Envelope vs Semantics
Table II displays the outcomes of the MM classification task conducted using two distinct stimulus features: speech envelope and word2vec features. The classification was done at the sentence level using data recorded under normal listening conditions. The results do not show a major difference in performance between the two features (Wilcoxon signed rank test, ). Therefore, it can be concluded that both acoustic and semantic features hold comparable importance in speech perception during normal listening conditions.
IV-C Dichotic Listening: Speech Envelope vs Semantics
The results, presented in Table II, demonstrate that the stimulus-response model using semantic features as input outperformed the model utilizing speech envelope features (Wilcoxon signed rank test, ). This indicates that semantic features are better entrained in the EEG response recorded during the dichotic listening experiment. The multi-modal network achieves an average accuracy of %, compared to % with the speech feature-based network. This significant performance improvement (Wilcoxon signed rank test, ) indicates that the EEG signal jointly encodes the semantic and acoustic content of the stimulus. Furthermore, this MM performance signifies that the proposed multi-modal model is able to correctly identify the intended channel of attention in more than % of the segments, a capability that can potentially influence BCI application design. The findings also suggest that semantics may hold higher significance than acoustic traits for speech perception under complex listening conditions.
IV-D Region-wise analysis
The region-wise analysis findings shown in Table VI, revealed that the temporal region exhibited the highest performance in the mismatch (MM) classification task, followed by the occipital, parietal, frontal, and central regions. The performance improvement is also observed to be consistent for semantic and acoustic features. This observed pattern suggests that certain brain regions may play more critical roles in processing dichotic listening stimuli, particularly in tasks involving attention and speech processing.
IV-E Attention to right ear versus left ear in dichotic listening
Additionally, the results presented in Table VII show that subjects exhibited significantly greater attention to the stimulus presented in the right ear (S2) compared to the left ear (S1). This supports the phenomenon of right ear dominance[43] in dichotic listening tasks, where individuals typically pay greater attention to stimuli delivered to the right ear.
V Conclusion
In this study, we compared EEG-stimulus models under both natural speech and dichotic listening conditions, aiming to assess the relative contribution of acoustic and semantic features in different listening scenarios. Our findings reveal that the significance of textual and acoustic content is similar during natural speech listening. However, in the dichotic speech listening task, EEG signals demonstrated a higher match-mismatch performance for textual data compared to the audio signal. This suggests a more pronounced encoding of higher-level semantic information over acoustic envelope information in dichotic listening condition.
The study also investigated the role of word boundary information in the MM task performance. It is seen that the word boundary information plays a pivotal role in speech perception during the dichotic listening task. Notably, not all boundaries seem to be critical as in the case of natural speech. Further, the study added further evidence to right ear advantage observed in dichotic listening through the match-mismatch classification experiments.